OpenStudy (anonymous):

if f(x)=-5x-11 then f^-1(x)= ? inverse functions confuse me please help

OpenStudy (anonymous):

can you solve \[-5x-11=7\]?

OpenStudy (anonymous):

for example i mean it takes two steps only

OpenStudy (anonymous):

\[x=-\frac{ -18 }{ ?5}\]

OpenStudy (anonymous):

well \[\frac{18}{5}\]yes how did you do it?

OpenStudy (anonymous):

I did the work and checked it with a calculator

OpenStudy (anonymous):

"the work" being what ?

OpenStudy (anonymous):

but is that all there is to it?

OpenStudy (anonymous):

two steps right? what were those two steps?

OpenStudy (anonymous):

switch x and y, solve for y and x, and substitute f^-1 for y

OpenStudy (anonymous):

my school teaches it in three steps

OpenStudy (anonymous):

but that isn't the answer they want

OpenStudy (anonymous):

ok all i was trying to point out is that if you can solve \[-5x-11=7\] that means if you know the output 7, you know the input \(\frac{18}{5}\) the two steps where a) add 11 b) divide by \(-5\) that is what the inverse does

OpenStudy (anonymous):

if you cannot do it in your head (which you can) solve \[-5y-11=x\] for \(y\) same two steps a) add 11\[-5y=x+11\]divide by \(-5\) \[y=\frac{x+11}{-5}\]

OpenStudy (anonymous):

or if you want to pull out the minus sign, write \[f^{-1}(x)=-\frac{x+11}{5}\]